OpenStudy (anonymous):

By showing all steps, solve for x if the perimeter of a triangle is 26. Side A: 14-2x Side B:3x-1 Side C: 2x+1 (I'll medal and fan)

OpenStudy (campbell_st):

well you know Perimeter = sum of the 3 sides so 26 = 14 - 2x + 3x -1 + 2x + 1 collecting like terms and you get 26 = 3x + 14 all you need to do now is solve for x hope it helps

OpenStudy (anonymous):

I'm lost, how did you get 26=3x+14?

OpenStudy (campbell_st):

well I collected the like terms when I added the 3 sides so looking at the x's -2x + 3x + 2x = (-2 + 3 + 2)x = 3x now the whole numbers 14 - 1 + 1 = 14 so the sum of the 3 sides of the perimeter is 3x + 14 and you are told it equals 26 so thats how you get the equation 26 = 3x + 14

OpenStudy (anonymous):

That helped me so much, thank you!